OpenStudy (anonymous):

Find the area and the circumference of a circle with radius 9cm . Use the value 3.14 for (pi) , and do not round your answers. Be sure to include the correct units in your answers.

OpenStudy (mathmagician):

C=2*pi*R=2*3.14*9cm=56.52cm

OpenStudy (anonymous):

is the circumfrance?

OpenStudy (anonymous):

Yes. Area = 3.14 x 9 * 9 = 254.34 units^2